NAOJ GW Elog Logbook 3.2

In the first picture, I showed the component on the optical bench now, and the red line represents the infrared and green line for green. As it shows in the picture, the transmit beam of the cavity divided by the BS mirror, the reflection part goes into the PD and the transmission part goes into the CCD, so I can see the shape of the beam through the screen connect to it.
And the second picture shows the loop I used to lock the cavity,the red line stands for the beam, I only list the optical component which connected to this loop. The black lines and the orange lines show the wire, but the wires in orange only used for the first step.
As I mentioned in my last report, the Function Generator on the left provides a 15.2MHz sin wave, which is sent to EOM to do the modulation, and also there is the other line come out from this, goes into the 'LO'(Local Oscillate) port of the demodulation board.
Then from the output of PD also goes to two different ways, one goes to the 'RF'( radio frequency) port of the demodulation board and the other goes to the channel 1 of the oscilloscope which shows in yellow on the screen of it.
The signal of 'RF' and 'LO' mixed together in the demodulation board and the output of it will go through the low pass filter of the board and then connected to the channel 2 of the oscilloscope, it shows the error signal and it is blue.
At the first beginning, I used the Function Generator on the right to send a triangle wave to the PZT driver to scan it and find the mode matching. Then when I got the error signal(the output of the demodulation board), firstly, this signal was sent to Stanford who acts as a low pass filter and then the output of Stanford goes to the PZT driver also.Now the loop has been closed.
And then I stopped scanning the cavity(plug out the orange lines), but just change the voltage of the PZT driver by myself. Until from the CCD screen, I am around the maximum of the TEM00 modes. Then I checked, the cavity is locked and the green beam is very powerful.
That's how I did until now, but actually the lock can not keep for very long, only last about 20-30 mins. So since I cannot switch off the loop(there is no switch I can use), I just need to plug out one off the wire and change the voltage of the PZT driver near the resonance and recover the loop. Try to do it this way and find out if I can get better result or not.




Before start to lock the cavity, I set the temperature to the phase matching temperature which we found quiet a while ago. The readjust the mirror a little to bit to get the best mode matching I can. Then I set the the sinusoidal wave with frequency of 15.2MHz and amplitude of 400mV, sent it to the EOM to do the modulation,and this is also used to be the 'LO' signal in the demodulation process. Then used the PD to get the transmitted beam of the cavity, which is the 'RF' signal of the demodulation. Then with the demodulation board, the 'LO' and 'RF' signal mixed together, the output( error signal ) sent to the oscilloscope.Also the error signal will go through the low pass filter and send to the input of the PZT driver. The other part of the PZT input is a triangle signal(parameter: 28Hz, amplitude: 2.5V, offset:2.5V) As soon as I got the error signal I stopped scan the cavity with triangle wave and adjust the PZT by changing the offset of PZT driver until the cavity is locked. After the cavity locked, the green light become much more powerful, and the reflection of the green light goes back into the Faraday Isolator, and the reflection from the Faraday point to the height around people's eyes which is really dangerous.

I temporarily moved another optical table inside the clean booth because I need some extra space in a clean environment to test the new assembly of the translation stage.

Today we clean up the unused part of the optical bench for the two lasers, and we found a rack for the power supply. The arrangement shows in the second picture.

In order to decrease the height of the sample on the translation stage, I designed a new configuration of the assembly of the translation stage and some new parts. With the help of Hirata san, I made these drawings and submitted at the mechanical shop of ATC.


I tidied up the power supply of the absorption measurement system because it was too messy. There were a lot of unnecessary multiple sockets connected in series (which is bad for the electrical grounding), especially under the stairs (all covered by dust). So, I rearranged the multiple sockets and put the unneeded ones on a shelf in the first floor, see the picture.




I upload the front panels and diagrams of the current version of the VIs and subVIs that I made with the help of Sakai-san.
The main VIs are "Translation Stage Map.vi" and "Translation Stage Scan.vi"
I don't upload the diagrams of subVIs that belong the SR830's and the Zaber motors' libraries.

L1: 37 cm
L2: 46 cm
L3: 55 cm
the final beam waist is 52.7 micrometer and the position is at 82.2 cm.
The first picture shows what we have on the optical bench before holiday, it is only for doing the measurement, so the dichroic mirror is not in the right gesture. So today I flipped the dichroic mirror, and put the SHG housing in the right position and direction, according to the result showed in the last paragraph. Aligned the inject beam and the reflection beam, got the transmission beam after the SHG and the reflection beam at the back side of the mirror after EOM. I tried to find some mode but did not get any. I will try to do it tomorrow.
Another problem is that I thought the stage we are using now is broken...in one of the direction, it can only move one way, but it can still be used.

This week we increased the height of the crystal instead of lower the miniscope. But after the increase, we still cannot get rid of the higher modes. We guessed it is because of the mismatch of the beam waist size. Since the increase of the crystal was reached with putting a spacer under the 'L' part which used to fix the crystal, the wire of the thermal sensor and the peltier are also connected to this part, we cannot push too much on the crystal,so now when we touch the wire outside the housing, the position of the crystal will change slightly. Another problem is that we found out the first turning mirror is too closed to the beam waist focused by the first lens of the telescope, and this caused the optical damage of the mirror coating, since the power density at the beam waist is too high when we use the full power of the laser. So we decided to move the the easily damaged turning mirror one hole closed the first lens, and do the same change with all the other component of the telescope without changing the distance between them. Then according to the new version optical scheme, we changed the dichroic mirror into the one transmit green and reflect red. In this case, we can use this dichroic mirror and the last mirror of the telescope to adjust the beam, things became easier. As I mentioned before, we guess the reason of the higher mode existence is because of the mismatch of the beam size. So we turned the cavity again,inject the beam from the back side of the housing, tried to lock the cavity and get the exact beam waist size we need. Give the EOM 15.235MHz modulation and use the PD to get the error signal. With the demodulation board found in TAMA, and the Stanford as a low pass filter, we locked the cavity, but cannot get rid of the offset. The lock last for more than two hours until we swtiched off everything and left. During the lock, we measure the beam size and got the beam waist size the cavity need is 54.1 micrometer. So the next step is the adjust the telescope again to get the beam waist size given by this locking.

Picomotors of the 4 suspended mirrors (2 of filter cavity and 2 of telescope) need to be put in operation.
The cables that connect the picomotor to the the chamber flange inside the chamber should be already in place for the 4 suspension (to be checked)
Takahashi-san provided driver for the picomotor (fig 1) to be connected to a cable (fig 2) to the extern part of the flange. This device were used to control picomotors of TAMA SAS.
The drivers can be controlled by a joystick and by remote using a labview software that Takahashi-san will provide
Some more information:
1- Picomotors drivers are 3 (IM1 IM2 NM3), NM2 could not be found. We checked in both end rooms without success.
2- We need 4 cable to connect them too the 4 chambers. For the moment we have only 2. (Again we checked in both the end rooms). According to Takahashi-san the missing two cable can be built using spare connectors like that in figure 2 and spare suitable cables I already found in TAMA. Takahashi-San can also provide the suitable tool to connect pins in the connector.
3- All this material is temporarily stored in a box under my desk in TAMA

In the past days with the valuable help of Matteo L. we realised 4 NIM-modules able to filter an input signal with a lowpass filter of 2 order (cut frequency at 100 Hz) and to amplify it of a factor 100.
The scheme of the circuit, using a multiple feedback (MFB) filter, drawn with the software LTSpace is shown picture 1 and its transfer function in picture 2.
Each module can process two channels, so in total we have 8 available filters+amplifiers.
A picture of the circuits and the module are shown in attached files 3 and 4.
They will be used to filter and amplify the output signal of the PSD used for the filter cavity optical levers, replacing the stanfords that have been doing this job up to now.



Today we restart the alignment from the first beginning. Firstly, with the power meter, measure the input and output of the EOM, adjust the transfer stage until almost the same power value. Then for the component, try to adjust them until the beam shine on the middle of everyone. With the tilt angle of the mirror, adjust the beam height as 76cm at both the near field and far field. Finally, we found out the transfer stage under the SHG housing is too high for present beam height.So we remove the bottom layer, got the transfer beam and the reflection beam, but the green is weak. The possible reason may be:1.Low input power 2.During the adjust process, the position of the lens may change,leading to the change of the waist position, so the input beam size of SHG crystal is too large 3.The polarization 4. The misalignment of the beam height and the crystal height. We will check all the possible reason tomorrow.

Last Friday, we installed the mirror part of the cavity,tried to align it. For doing this, firstly we add 20 kiloohm resistance to the present photo diode to increase the sensitivity at the cost of reducing the bandwidth(200 kHz). At the first try,we got multiple modes.With two mirrors and the lens, we tried to rule out higher modes and get the TEM00 one, but found out that there is misalignment in the vertical direction. After adding some piece under the mirror block, this problem has been solved, but still we cannot get rid of the effect of TEM02(TEM20) mode. We will restart the alignment again today, try to make some progress.